# Democratizing Large Language Models via Personalized Parameter-Efficient Fine-tuning

### v1

@@ 4.8. On PEFT Method Choices
We envision the proposed OPPU as a versatile, PEFT-based LLM personalization framework compatible with various PEFT methods. This study demonstrates OPPU’s performance across different PEFT approaches. Beyond the most commonly adopted LoRA, we explore prompt tuning and $\mathrm{(IA)^{3}}$, which plug in external learnable parameters in the embedding space and scale the attention factor, respectively. As Figure 7 illustrates, the OPPU framework enhances performance with all three PEFT types, demonstrating the effectiveness and versatility of OPPU across various PEFT methods. Notably, LoRA typically delivers the highest performance, followed by $\mathrm{(IA)^{3}}$, and then prompt tuning. This hierarchy aligns with the proportion of trainable parameters in each method: LoRA at 0.01%, $\mathrm{(IA)^{3}}$ at 0.06%, and prompt tuning at 0.001%. These results suggest that a greater number of trainable parameters in a personalized PEFT method generally leads to improved personalization performance.
4.9. Case Study
To provide a qualitative understanding of OPPU, we conduct a case study focusing on an individual user in the personalized movie tagging task. As shown in Figure 6, the non-personalized method overlooks the user’s behavior history and bases its prediction solely on the user’s query input, leading to an evidently incorrect answer. The retrieval-based method incorporates user behavior history by retrieving the most relevant user history items, but it falls short of identifying the user’s most pertinent historical behaviors that mirror the query, leading to incorrect LLM output. We argue that retrieval augmentation only involves user history data via several retrieval examples, limiting the comprehensive understanding of user preferences. In contrast, our proposed OPPU employs a personalized PEFT module to effectively capture the user’s behavior preferences and patterns across the entire user history corpus. In this case, based on the user’s most frequently tags movies ”based on a book” in behavior history, OPPU successfully identifies this pattern and consequently provides the correct response.
5. Conclusion
@@
LLMs personalization has emerged as a rapidly evolving research area, aiming to tailor LLMs’ emergent abilities to users’ unique needs. In this work, we introduce OPPU, which utilizes personalized PEFT parameters as a proxy for LLM personalization, demonstrating the advantage of model ownership and enhanced generalization under user behavior shift. By tuning these parameters with a user’s behavioral history, OPPU encapsulates the user’s behavioral history and patterns in private PEFT parameters. Furthermore, OPPU can be integrated with non-parametric user knowledge via history retrieval and user profile augmentation, exhibiting state-of-the-art performance across all seven public tasks in the LaMP benchmark. Additional experiments highlight OPPU’s versatility, robustness, and superiority in modeling users across different active levels. Our proposed OPPU framework paves the way for new opportunities in PEFT-based LLM personalization, enhancing the modularity of LLMs for more effective and democratizing personalization.
Appendix A Limitations
@@
We identify two key limitations in OPPU. Firstly, limited by the dataset, we mainly focus on one specific task per user rather than examining user behaviors across multiple tasks and domains. For example, in the movie tagging task, users are solely engaged in that specific activity, without the inclusion of behaviors from other areas. Despite this, the OPPU framework is inherently adaptable to any text sequence generation task and is capable of conducting diverse user instructions across different tasks and domains. The exploration of LLM personalization across a broader range of tasks and domains remains an area for future investigation. Secondly, OPPU serves as a general framework that incorporates the entirety of a user’s behavior history into their private PEFT module. However, user interests are dynamic and may display inconsistencies or conflicts over time. Future research directions include examining methodologies for selecting the most relevant or valuable items from a user’s history and devising strategies to effectively manage any discrepancies or conflicts within this historical data.
Table 5. Hyperparameter settings of OPPU accross various tasks on LaMP benchmark. We find our hyperparameter settings robust across all 7 tasks.
Tasks rank #epoch lr R2 reg. batch size
LaMP-1: Personalized Citation Identification 8 3 $1e^{-5}$ $1e^{-2}$ 16
LaMP-2: Personalized News Categorization 8 3 $1e^{-5}$ $1e^{-2}$ 16
LaMP-2: Personalized Movie Tagging 8 3 $1e^{-5}$ $1e^{-2}$ 4
LaMP-3: Personalized Product Rating 8 3 $1e^{-5}$ $1e^{-2}$ 3
LaMP-4: Personalized News Headline Generation 8 2 $1e^{-5}$ $1e^{-1}$ 8
LaMP-5: Personalized Scholarly Title Generation 8 2 $1e^{-5}$ $1e^{-1}$ 4
LaMP-7: Personalized Tweet Paraphrasing 8 2 $1e^{-5}$ $1e^{-1}$ 8
Appendix B Ethics Statement
@@
Appendix C Hyperparameters
The hyperparameters of OPPU are presented in Table 5 to facilitate further research.
Figure 8. Efficiency analysis of OPPU, in which we alter the number of history items and average token per history item and record the training time.
Appendix D Efficiency Analysis
@@
Appendix E Scientific Artifacts
OPPU is built with the help of many existing scientific artifacts, including PyTorch (Paszke et al., 2019), Numpy (Harris et al., 2020), huggingface transformers (Wolf et al., 2020), and bitsandbytes (Dettmers et al., 2022). We will make the OPPU implementation publicly available to facilitate further research.
Appendix F Computation Resources Details
@@
Appendix G PEFT Cosine Similarity Details
@@
Appendix H Task Details
@@
Appendix I Prompt Details
@@
I.1. Personalized Citation Identification
@@
I.2. Personalized News Categorization
@@
I.3. Personalized Movie Tagging
@@
I.4. Personalized Product Rating
@@
I.5. Personalized News Headline Generation
@@
I.6. Personalized Scholarly Title Generation
@@
I.7. Personalized Tweet Paraphrasing

### v2

@@ 6 Analysis
We propose OPPU as a versatile PEFT-based LLM personalization framework compatible with various PEFT methods. This study evaluates OPPU’s performance across different PEFT approaches, including LoRA, prompt tuning, and $\mathrm{(IA)^{3}}$, which plug in external learnable parameters in the embedding space and scale the attention factor, respectively. As shown in Figure 6, OPPU enhances performance with all three PEFT types, demonstrating its effectiveness and versatility. Notably, LoRA typically delivers the highest performance, followed by $\mathrm{(IA)^{3}}$, and then prompt tuning. This hierarchy aligns with the proportion of trainable parameters in each method: LoRA at 0.01%, $\mathrm{(IA)^{3}}$ at 0.06%, and prompt tuning at 0.001%. These results suggest that a greater number of trainable parameters in a personalized PEFT method generally leads to improved personalization performance.
Case Study
To illustrate the effectiveness of OPPU, we conduct a case study on personalized movie tagging task for an individual user. Figure 5 shows that the non-personalized method, relying solely on query input, ignores user behavior history and yields incorrect answers. The retrieval-based method, though incorporating user history, fails to retrieve closely matched behaviors to the query, also resulting in errors. We argue that retrieval augmentation with a few user history examples cannot fully capture user preferences. In contrast, OPPU uses a personalized PEFT module to effectively understand the user’s behavior patterns across the entire user history. In this case, OPPU successfully recognizes the user’s frequent tagging of “based on a book" and provides the correct response.
Figure 7: Similarities between personal PEFT parameters under personalized text classification and generation.
Similarities Between Personalized PEFTs
To understand how user behavior patterns are reflected in their private PEFT parameters, we analyze the cosine similarities between these parameters across different users, as shown in Figure 7. We select two representative tasks from text classification and generation categories and compute the cosine similarities for 100 users’ PEFT parameters in the test set. The private PEFT similarities generally range from 0.4 to 0.7, with the highest average similarities observed in the scholarly title generation task, likely due to its task-specific nature. Relative differences among users offer additional insights: in personalized text classification tasks, similarities vary more, indicating that some users have higher similarities than others. Conversely, in personalized text generation tasks, the similarities are relatively uniform, suggesting that personal preferences in these tasks are harder to categorize.
7 Related Work
7.1 Personalization of LLMs
The thrust of existing LLM personalization research is centered on designing prompts that incorporate historical user-generated content and behavior. These approaches help LLMs understand users’ preferences, tailoring responses to individual needs Tan and Jiang (2023); Chen et al. (2023). The endeavors towards personalized LLMs mainly fall into three categories: vanilla, retrieval-augmented, and profile-augmented personalized prompts.
In the vanilla personalized prompt category, researchers use in-context and few-shot learning to encode either complete or a sample of user behavior history as contextual examples Liu et al. (2023a); Wang et al. (2023). For instance, Dai et al. (2023) and Kang et al. (2023) encode the user’s personal rating history as few-shot demonstration examples. Moreover, some research works Christakopoulou et al. (2023); Zhiyuli et al. (2023) also discovered a long user history would bring better performance. To manage the growing user behavior data and LLMs’ limited context windows, the retrieval-augmented personalized prompt approach has emerged Salemi et al. (2023); Li et al. (2023). For instance, Pearl Mysore et al. (2023) proposes a generation-calibrated retriever to select historic user-authored documents for prompt augmentation. Beyond simple retrieval, some researchers summarize user preferences and behavior patterns into natural language profiles for input query augmentation, termed profile-augmented personalized prompts Liu et al. (2023b); Sun et al. (2024). Richardson et al. (2023) use the instruction-tuned LLMs to generate an abstract summary of user history data, augmenting retrieval-based personalization methods. There is also another line of work focusing on personalized alignment methods via parameter merging Jang et al. (2023) and personalized reward model Cheng et al. (2023).
7.2 Parameter-Efficient Fine-tuning (PEFT)
With the exponential growth in LLM parameters, fine-tuning all parameters is expensive Liu et al. (2022b); Xu et al. (2023); Gupta et al. (2024); Jian et al. (2024). To address this, parameter-efficient fine-tuning (PEFT) methods update only a small number of extra parameters while keeping pretrained weights frozen He et al. (2021); Fu et al. (2023); Liu et al. (2024); Dou et al. (2024); Zhang et al. (2024). For example, adapter tuning Houlsby et al. (2019) injects learnable parameters into each feedforward layer, updating only these during fine-tuning. Inspired by discrete textual prompts Sanh et al. (2022); Wang et al. (2022), prefix tuning Li and Liang (2021) and prompt tuning Lester et al. (2021) optimize prompts and prefixes for specific tasks. LoRA Hu et al. (2021) adds low-rank matrices to approximate parameter updates, and $\rm{(IA)^{3}}$ Liu et al. (2022a) scales activation in the attention mechanism. These methods achieve performance comparable to full fine-tuning by updating less than 1% of the original parameters, are effective against catastrophic forgetting Pfeiffer et al. (2021), and are robust to out-of-distribution samples Li and Liang (2021).
Previous works focused on prompt design, limited by model ownership and user behavior shifts. PEFT’s small number of updated parameters and plug-and-play nature make it ideal for efficient LLM personalization and model ownership. OPPU introduces personalization at the parametric level via a personal PEFT module, pioneers storing user history within personal PEFT parameters, equipping each user with a unique, easily integrable PEFT module for model ownership.
8 Conclusion
@@
We introduced OPPU, equipping each user with a personal PEFT module that facilitates model ownership and generalization under behavior shifts. By tuning these parameters with a user’s history, OPPU captured personalized behavioral patterns. It integrated non-parametric user knowledge via retrieval and user profiles, showing superior performance across all seven LaMP benchmark tasks. Additional experiments demonstrated OPPU’s versatility, robustness, and effectiveness for users with varying activity levels. Our framework paved the way for new opportunities in PEFT-based LLM personalization, enhancing LLM modularity for effective and democratized personalization.
9 Limitations
@@
We identify three key limitations in OPPU. Firstly, limited by the dataset, we mainly focus on one specific task per user rather than examining user behaviors across multiple tasks and domains. For example, in the movie tagging task, users are solely engaged in that specific activity, without the inclusion of behaviors from other areas. Despite this, the OPPU framework is inherently adaptable to any text sequence generation task and is capable of conducting diverse user instructions across different tasks and domains. The exploration of LLM personalization across a broader range of tasks and domains remains an area for future investigation. Secondly, OPPU serves as a general framework that incorporates the entirety of a user’s behavior history into their private PEFT module. However, user interests are dynamic and may display inconsistencies or conflicts over time. Future research directions include examining methodologies for selecting the most relevant or valuable items from a user’s history and devising strategies to effectively manage any discrepancies or conflicts within this historical data.
10 Ethical Considerations
@@
11 Acknowledgements
This work was supported by NSF IIS-2119531, IIS-2137396, IIS-2142827, IIS-2234058, CCF-1901059, and ONR N00014-22-1-2507.
Appendix A Hyperparameters
The hyperparameters of OPPU are presented in Table 5 to facilitate further research. For LoRA, we add trainable low-rank matrice in the $W_{q}$ and $W_{v}$
Figure 8: Efficiency analysis of OPPU, in which we alter the number of history items and average token per history item and record the training time.
Appendix B Efficiency Analysis
@@
Appendix C Baseline Details
The baseline details are presented as follows:
  • Retrieval-Augmented Personalization (RAG): We follow the retrieval-augmented personalization method presented in LaMP Salemi et al. (2023), where the user’s query is augmented with top $k$ retrieved items from the corresponding user’s history corpus. We take $k$=1, 2, 4 in this work.
  • Profile-Augmented Personalization (PAG): This method is taken from Richardson et al. (2023), in which the user’s input sequence would concatenate the user’s profile summarizing the user’s preference and behavior patterns. In our experiments, we generate user profiles using the vicuna-7B Chiang et al. (2023) model. Moreover, the profile-augmented method could be combined with the retrieval augmentation. In this case, we take the number of retrieval items $k$=1 following the setting of Richardson et al. (2023).
Appendix D Scientific Artifacts
OPPU is built with the help of many existing scientific artifacts, including PyTorch Paszke et al. (2019), Numpy Harris et al. (2020), huggingface transformers Wolf et al. (2020), and bitsandbytes Dettmers et al. (2022). We will make the OPPU implementation publicly available to facilitate further research.
Appendix E Computation Resources Details
@@
Appendix F PEFT Cosine Similarity Details
@@
Appendix G Task Details
@@
Appendix H Prompt for Personalization Tasks
@@
H.1 Personalized Citation Identification
@@
H.2 Personalized News Categorization
@@
H.3 Personalized Movie Tagging
@@
H.4 Personalized Product Rating
@@
H.5 Personalized News Headline Generation
@@
H.6 Personalized Scholarly Title Generation
@@
H.7 Personalized Tweet Paraphrasing
@@
Appendix I Prompt for User Profile Generation
For user profile generation, we follow the prompt template in Richardson et al. (2023).
I.1 Personalized Citation Identification
Write a summary, in English, of the research interests and topics of a researcher who has published the following papers. Only generate the summary, no other text. User History: {USER HISTORY} Answer:
I.2 Personalized News Categorization
Look at the following past articles this journalist has written and determine the most popular category they write in. Answer in the following form: most popular category: <category>. User History: {USER HISTORY} Answer:
I.3 Personalized Movie Tagging
Look at the following past movies this user has watched and determine the most popular tag they labeled. Answer in the following form: most popular tag: <tag>. User History: {USER HISTORY} Answer:
I.4 Personalized Product Rating
Based on this user’s past reviews, what are the most common scores they give for positive and negative reviews? Answer in the following form: most common positive score: <most common positive score>, most common negative score: <most common negative score>. User History: Answer:Look at the following past movies this user has watched and determine the most popular tag they labeled. Answer in the following form: most popular tag: <tag>. User History: {USER HISTORY} Answer:
I.5 Personalized News Headline Generation
Given this author’s previous articles, try to describe a template for their headlines. I want to be able to accurately predict the headline gives one of their articles. Be specific about their style and wording, don’t tell me anything generic. User History: {USER HISTORY} Answer:
I.6 Personalized Scholarly Title Generation
Given this author’s previous publications, try to describe a template for their titles. I want to be able to accurately predict the title of one of the papers from the abstract. Only generate the template description, nothing else. User History: {USER HISTORY} Answer:
I.7 Personalized Tweet Paraphrasing
Given this person’s previous tweets, try to describe a template for their tweets. I want to take a generic sentence and rephrase it to sound like one of their tweets, with the same style/punctuation/capitalization/wording/tone/etc. as them. Only give me the template description, nothing else. User History: {USER HISTORY} Answer: